OpenStudy (anonymous):

Short Answer: Raise the quantity in parentheses to the indicated exponent, and simplify the resulting expression. Express answers with positive exponents. (28xy^3/130x^3y^-2)^3

OpenStudy (amistre64):

its best to write this by splitting apart the fraction there in order: 28 x y^3 ---- --- - --- 130 x^3 y^-2 its just easier to work off of then

OpenStudy (amistre64):

28 14 ---- = --- = 2/9 130 65 x x 1 --- = --- = --- = 1/x^2 x^3 xxx xx y^3 yyy ---- = ------- = yyyyy = y^5 y^-2 1/y 1/y 2y^5 ---- ^3 x^2

OpenStudy (amistre64):

2^3 y^5^3 8 y^15 ---------- = ------ x^2^3 x^6

OpenStudy (amistre64):

14/65 aint 2/9 is it ....

OpenStudy (amistre64):

9.7 = 63, doh!!

OpenStudy (amistre64):

well, the number parts are off, but its the same concept :)

OpenStudy (amistre64):

14^3 ---- = whatever the calculator gets us ;) 65^3

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

Will do the math in a sec gonna eat lunch

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

2744/274625